15 men can do a work in 210 days.if 15 additional men engaged every after 10 days, in how many days the work be finished?
Answer by ankor@dixie-net.com(22740)

15 men can do a work in 210 days.
Find how many man-days required to complete the job
15 * 210 = 3150 man days
:
if 15 additional men engaged every after 10 days, in how many days the work be finished?
:
Days - manhr - Tot manhr
10 - 15*10=150 - 150
20 - 30*10=300 - 450
30 - 45*10=450 - 900
40 - 60*10=600 - 1500
50 - 75*10=750 - 2250
60 - 90*10=900 - 3150 man-hrs complete the job
:
It will take 60 days and a total of 90 men will be working then